A TRANSITION SCHEME FOR OBSERVING SODIUM DIMER RYDBERG STATES AND PRODUCING STATE-SELECTED ULTRACOLD MOLECULAR IONS BY ALL-OPTICAL TRIPLE RESONANCE SPEC-TROSCOPY

\underline{J. LI}, J. ZHANG, J. BAHNS, W. C. STWALLEY, Department of Physics, University of Connecticut, Storrs, CT 06268; E. S. CHANG, Department of Physics & Astronomy, University of Massachusetts, Amherst, MA 01003; and C. C. TSAI, Department of Physics, National Chung-Hsing Univerity, Taichung, Taiwan, R.O.C.

The X^{1}Σ_{g}^{+}(1, J′′′) ⟶ A^{1}Σ_{u}^{+}(1, J′′) ⟶ 3^{1}Σ_{g}^{+}(0, J′) ⟶ nl^{1}Λ_{u}^{+}(0, J) transition scheme is used to study high n sodium dimer rydberg states. The spectrum is simplest when the intermediate state has J′ = 0 which implies J = 1. Analysis of our data yield an ionization potential of 39478.13±0.03cm^{−1}, implying a molecular ion dissociation energy of D_{0}^{0}(Na_{2}^{+})=7914.19±0.03cm^{−1}. The same transition scheme can be used to detect ultracold ground state molecules Na_{2}(v = 0, J = 0) and to produce ultracold ground state molecule ions Na_{2}^{+}(v^{+} = 0, N^{+} = 0).